OpenStudy (benlindquist):

If S is 20 when T is 4, then T is when S is 30.

OpenStudy (benlindquist):

is it 6

OpenStudy (welshfella):

30/20 = x /4 x = ( 4*30) / 20 = 6 yes

OpenStudy (benlindquist):

ok